OpenStudy (anonymous):

The circle below is centered at the point (2, -1) and has a radius of length 3. What is its equation?

Vocaloid (vocaloid):

the equation for a circle is (x-h)^2 + (y-k)^2 = r^2 now we plug in h = 2, k = -1, and r = 3

OpenStudy (anonymous):

I think its (x - 2)2 + (y + 1)2 = 32

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

:D